OpenStudy (blackstreet23):

(sin theta + cos theta)^2 =1+ sin 2 theta? how do i very this is an identity?

OpenStudy (anonymous):

square and you will see it

OpenStudy (campbell_st):

this is reasonably straight forward... you need to expand a perfect square (a + b)^2 = a^2 + 2ab + b^2 so what would your problem look like..?

OpenStudy (blackstreet23):

Sin^2 + 2 cos * sin + cos ^2

OpenStudy (blackstreet23):

And then?

OpenStudy (campbell_st):

great so what if I write it this way \[\sin^2(x) + \cos^2(x) + 2\sin(x)\cos(x)\] what do you know about \[\sin^2(x) + \cos^(x) ~~~~and~~~~~2\sin(x)\cos(x)\]

OpenStudy (campbell_st):

oops \[\sin^2(x) + \cos^2(x)\]

OpenStudy (blackstreet23):

Saw it thanks !!!